A student walks and jogs to school every day. She averages 4km/hour walking and 8km/hour jogging. The distance from home to school is 6km and the trip takes one hour. How far does the student walk.
Answer by mananth(12270)

walk = 4kph
jog 8 kph
..
distance to school = 6 km.
let her walk x km
she will jog 6-x km
..
Time walking = x/4
Time jogging = 6-x / 8
...
Total time = 1 hour
x/4 + (6-x)/8 =1
LCD = 8
(2x+6-x) / 8 =1
x+6 = 8
x= 2 km walking distance